### v1

# Nudged-Elastic Band Calculations of Polymorph Transitions and Solid-State Reactions in Molecular Crystals

Abstract
The modeling of solid-state transformations, such as polymorphic transitions and chemical reactions in molecular crystals, is vital for many applications like drug design or the development of new synthesis methods but faces several crucial challenges. First, the automatic initial pathway generation in terms of nudged-elastic band (NEB) calculations based on a linear interpolation often fails for periodic systems, leading to unrealistic geometries and atomic collisions. Second, the necessary system sizes are typically beyond the scope of density functional theory (DFT) calculations in terms of computational cost, but the associated accuracy is vitally needed. To address these issues, we introduce a novel hybrid interpolation method that combines linear interpolation for cell parameters with spherical linear interpolation (SLERP) for molecular structures or intramolecular fragments, ensuring smooth and realistic transitions. In addition, we also benchmark the accuracy of the recently introduced SO3krates machine-learned force field (MLFF) architecture in this context, which has the potential to substantially improve the computational efficiency. We apply these methods to two polymorph transitions and a solid-state Diels-Alder reaction and show the applicability of our interpolation method. Furthermore, we find excellent agreement of MLFF results with DFT energy barriers and pathways, when trained on appropriate reference data, potentially paving the way for more automated explorations of solid-state transformations in molecular crystals.
I Introduction
Understanding solid-state transformations like transitions between molecular crystal polymorphs or solid-state chemical reactions at a molecular level is vital for pharmaceuticals, chemistry, and materials scienceKaupp (2002); Datta and Grant (2004); Anthony (2006); Saha, Nath, and Thakuria (2022). Polymorphs are different crystal-packing arrangements of the same molecule and although the chemical composition of a molecular crystal does not change when one polymorph transforms into another, this can often be associated with quite drastic changes in several properties. One important example crucial for the pharmaceutical industry is that polymorph transitions can affect the efficacy of drugsZhou et al. (2018). Chemical reactions in the solid state, as opposed to synthesis in solution, can open up novel reaction pathways and selectivitiesGalwey (2014). However, characterizing these transformations is challenging due to the complex interplay between molecular reorganization and the constraints imposed by the crystal lattice.
The primary computational approach to investigating reactions or transformations involves identifying the minimum-energy pathway (or reaction coordinate) connecting the initial and final states of a transformation. This is typically achieved through the use of the nudged elastic band (NEB) methodJónsson, Mills, and Jacobsen (1998); Henkelman, Uberuaga, and Jónsson (2000); Sheppard et al. (2012); Mader et al. (2018); Shin et al. (2019); Han and Luber (2020); Hossain et al. (2020); Podewitz, Sen, and Buchmeiser (2021), which efficiently samples the minimum-energy path by connecting a series of images or configurations along the reaction path. NEB is particularly effective for determining transition states and energy barriers, as it maintains a series of replicas of the system that are connected by springs, ensuring smooth transitions along the reaction coordinate. Therefore, this constitutes the main methodology for studying chemical reactions in vacuum or within implicit solvent models.
However, NEB calculations for solid-state systems are often hindered by the complexity of generating realistic initial reaction pathways, particularly for periodic molecular crystal structures. Linear interpolation, the simplest method for generating these paths, often fails to capture the intricate changes in both molecular orientation and unit cell parameters that occur during solid-state transformations, frequently leading to atomic collisions or nonphysical geometries. While in isolated systems problems with the interpolation can be corrected by hand, such manual corrections are extremely cumbersome for complex periodic structures.
Another promising method for modeling dynamic processes is molecular dynamics (MD), which has been applied to polymorphic transitions of molecular crystals utilizing traditional force fieldsSalvalaglio et al. (2014); Gimondi and Salvalaglio (2017); Francia, Price, and Salvalaglio (2021). However, most traditional force fields are not designed to handle the breaking of chemical bonds, which needs to occur in chemical reactions, and their accuracy is often not sufficient. The main workhorse for the modeling of molecular crystalsOtero-de-la-Roza and Johnson (2012); Reilly and Tkatchenko (2013); Marom et al. (2013); Reilly and Tkatchenko (2015); Brandenburg and Grimme (2016); Shtukenberg et al. (2017); Hoja et al. (2019); Dolgonos, Hoja, and Boese (2019); Jana et al. (2021); Price, Otero-de-la-Roza, and Johnson (2023); Hoja, List, and Boese (2023) is density functional theory (DFT), since it provides a good balance between accuracy and computational cost. While first-principles MD simulations can be performed for small molecules in vacuum, modeling the kinetics of polymorph transitions and solid-state reactions with either MD or NEB for realistic systems is typically out of reach in terms of available computation times. In recent years, huge advancements have been made in the field of machine-learned force fields (MLFF), especially in the development of neural networksSchütt et al. (2017); Schütt, Unke, and Gastegger (2021); Unke et al. (2021); Batzner et al. (2022); Batatia et al. (2022); Frank, Unke, and Müller (2022); Frank et al. (2024). With an appropriate set of accurate reference data — typically high-level DFT calculations — MLFFs can be trained so that they are almost as efficient as traditional force fields but also much more accurate for specific applications.
In this paper, we present a novel interpolation method to generate high-quality initial reaction paths for NEB calculations. The proposed method is designed to handle both rigid and flexible molecular systems, automatically avoiding atomic collisions and ensuring smooth transitions between the initial and final configuration. Furthermore, we also benchmark the applicability and accuracy of state-of-the-art MLFFs in the context of periodic NEB calculations. For that we generate tailor-made DFT reference data and utilized the newly developed equivariant message-passing neural network architecture SO3kratesFrank, Unke, and Müller (2022); Frank et al. (2024).
This methodology is then tested on the polymorph transitions of the (low-temperature) polymorph of 1,4-diiodobenzeneAlcobé et al. (1994) into the (high-temperature) polymorph as well as that of theophylline polymorph I Khamar et al. (2011) into polymorph II.Fucke et al. (2012a) In addition, we have also modeled the solid-state reaction of a cocrystal composed of bis(N-allylimino)-1,4-dithiin and 9-bromoanthracene into the Diels-Alder productKhorasani et al. (2015). We show that our new interpolation method automatically produces viable initial reaction paths and that the obtained MLFFs are providing very accurate results compared to the much more expensive DFT reference data.
II Computational Methods
For the three considered transitions, we have utilized the following experimentally determined crystal structures as starting point, with the Cambridge structural database code listed in parenthesis: 1,4-diiodobenzene polymorph (ZZZPRO05), 1,4-diiodobenzene polymorph (ZZZPRO07), theophylline polymorph I (BAPLOT04), theophylline polymorph II (BAPLOT06), co-crystal of bis(N-allylimino)-1,4-dithiin and 9-bromoanthracene (FUSCIH), and the respective Diels-Alder product (FURROB). All these structures are visualized in Fig. 1.
First, all structures were fully optimized with PBE+MBD-NLPerdew, Burke, and Ernzerhof (1996); Tkatchenko et al. (2012); Ambrosetti et al. (2014); Hermann and Tkatchenko (2020) utilizing the 2020 light species default settings for basis functions and integration grids within FHI-aimsBlum et al. (2009); Knuth et al. (2015); Ren et al. (2012); Yu et al. (2018); Havu et al. (2009); Ihrig et al. (2015) (version 231212). The k-point grids for the DFT part and the MBD part were chosen so that the number of k-points Å or 30 Å, respectively, where is the cell length in the respective direction. We have applied the following convergence criteria: eV for the total energy, 10-5 electrons per Å3 for the charge density, 10-4 eV/Å for forces, 10-3 eV for the sum of the eigenvalues, and 0.005 eV/Å during geometry optimizations. All calculations were carried out via the Atomic Simulation Environment ASELarsen et al. (2017).
Since a simple linear interpolation does not allow the generation of reasonable starting structures for a reaction path sampling via nudged elastic band (NEB) calculations for all investigated systems, we have developed the following interpolation methodology. Our approach to NEB pathway generation introduces a hybrid interpolation method which uses a linear interpolation for the unit cell and leverages the strengths of the spherical linear interpolation algorithmShoemake (1985) (SLERP) for molecular structures
with and being the quaternions representing the initial and final structures, and being the interpolation parameter between 0 and 1. SLERP was for instance already utilized for protein calculationsNguyen, Jaillet, and Redon (2017, 2018).
To properly capture the changes of molecular structures during the transformation, it is necessary to specify bonds that are likely to rotate during the reaction. Therefore, we have defined intramolecular fragments for all molecules except 1,4-diiodobenzene. The fragment definitions are visualized in Fig. 2.
By dividing molecules into fragments, the SLERP algorithm is applied to each one of them individually and the bond lengths between the fragments are interpolated linearly, ensuring accurate representation of molecular flexibility as well as changes in the bond order. Furthermore, the overall molecular structure is shifted so that the center of mass is translated linearly within the unit cell across the pathway.
Since these fragment interpolations can lead to atomic collision in the unit cell for some images, we employ a post-processing step that adjusts the unit cell size to ensure chemically reasonable structures throughout the pathway. Therefore, we first detect possible atomic collisions by searching across all images for intermolecular distances smaller than times the respective covalent radius as used by ASE. In the image with the maximum number of collisions we obtain the smallest occurring intermolecular distance. Then, the unit cell vectors for that particular image are increased by the the respective projection of the identified distance vector. Then the linear interpolation of the unit cell vectors is repeated for the whole pathway utilizing the newly changed cell parameters as fixed intermediate image. This process is repeated until the resulting pathway is completely free of collisions.
This initial interpolation methodology was utilized throughout. For 1,4-diiodobenzene we have additionally also used a simple linear interpolation for comparison, for the other systems this leads to atomic collisions.
Next, we have performed nudged elastic band (NEB) calculations utilizing the solid-state NEBSheppard et al. (2012) as implemented in TSASE, which allows for variable unit cells during the reaction path sampling. We always applied the climbing image methodHenkelman, Uberuaga, and Jónsson (2000) and used as convergence criteria 0.005 eV/Å for 1,4-diiodobenzene and 0.1 eV/Å for the two remaining, more complex systems.
For 1,4-diiodobenzene we have performed NEB calculations with PBE+MBD utilizing the same settings as for the lattice relaxations described above. In this case, we have also performed a harmonic phonon calculation for the obtained transition-state estimate to validate the NEB results. For this calculation, we applied PBE+MBD utilizing phonopyTogo and Tanaka (2015) together with FHI-aims and a supercell of at least 12 Å was used to avoid artifacts. Furthermore, we have also performed a NEB pathway optimization for 1,4-diiodobenzene using DFTB3+MBD as implemented in DFTB+Hourahine et al. (2020), with a k-point grid satisfying Å.
Next, we have created PBE+MBD reference data for the training of machine-learned force fields (MLFF) suitable for NEB calculations. Therefore, we have performed for all systems 5 ps NPT molecular dynamics (MD) simulations at 500 K utilizing FHI-aims together with ASE. We have used a timestep of 1 fs, a characteristic timescale of 15 fs for the thermostat, and 2000 fs2 for the barostat constant (pfactor). These simulations were performed at ambient pressure as well as at 1 GPa. For 1,4-diiodobenzene and theophylline we have utilized a 1x2x2 and 1x3x1 supercell, respectively. In addition, we have performed full lattice relaxations of all initial reaction-path images directly after interpolation. In this way, the reference energies and forces should contain a sufficiently large portion of the potential energy surface, suitable for the specific NEB calculations.
The obtained PBE+MBD data was now used to train SO3kratesFrank, Unke, and Müller (2022); Frank et al. (2024) MLFFs for 2000 epochs, utilizing mostly default settings. For 1,4-diiodobenzene and theophylline we utilized 5000 samples for training and for the solid-state reaction 4000, while 1000 samples were always used for validation. Subsequently, lattice relaxations and subsequent NEB calculations were carried out for all systems in the same fashion as described for the PBE+MBD calculations above. We have utilized the GLPLanger, Frank, and Knoop (2023) package to obtain ASE calculators of our three MLFFs.
Finally, we have performed several single-point calculations on top of the obtained optimized NEB reaction paths. On top of the SO3krates structures, we have performed PBE+MBD/light single points within FHI-aims as described above. For the PBE+MBD/light reaction path of 1,4-diiodbenzene, we have additionally also calculated PBE+MBD energies utilizing tight species default settings and ME3(PBE0+MBD:PBE+MBD)/tight multimer embeddingHoja, List, and Boese (2023) energies, essentially approximating tight PBE0+MBD results.
III Results and Discussion
III.1 1,4-Diiodobenzene Polymorph Transition
We start by discussing the smallest considered system — 1,4-diiodobenzene. The low-temperature polymorph and the high-temperature polymorph differ mainly in the relative orientation of two of the four molecules within the unit cell. The remaining two molecules maintain a similar position and orientation in both crystalline structures. Experimentally, the transition temperature from the to the polymorph amounts to 326 K Alcobé et al. (1994), suggesting a quite small energy barrier for this transition.
Fig. 3 shows the obtained NEB results for several methods for the (left) to the (right) transition. The corresponding energy barriers (calculated from polymorph to the maximal energy) and the thermodynamic relative stability of the polymorphs , where a negative value indicates that polymorph is more stable, are listed in Table 1. The PBE+MBD/light results yield a small energy barrier of only 6.3 kJ/mol, which fits well with a transition close to room temperature. In terms of thermodynamics, both polymorphs have virtually the same stability, since we cannot properly distinguish stability differences of less than 1 kJ/mol. Please note, that the values reported here are relative static lattice energies (normalized per molecular unit), which do not include vibrational free energy effects. When performing tight PBE+MBD single points or tight PBE0+MBD estimates via multimer embeddingHoja, List, and Boese (2023) including up to trimer interactions (ME3), the relative energies change less than 1 kJ/mol.
| Energy Method | Opt. Method | ||
| PBE+MBD/light | PBE+MBD/light | 6.3 | -0.2 |
| PBE+MBD/tight | PBE+MBD/light | 5.4 | -0.4 |
| ME3/tighta | PBE+MBD/light | 6.4 | 0.0 |
| DFTB3+MBD | DFTB3+MBD | 4.0 | 3.3 |
| PBE+MBD/light | DFTB3+MBD | 7.0 | -0.6 |
| SO3krates | SO3krates | 6.0 | -0.4 |
| PBE+MBD/light | SO3krates | 6.5 | -0.5 |
a ME3(PBE0+MBD:PBE+MBD)/tight
When analyzing the NEB results obtained with the semi-empirical density functional tight binding approach DFTB3+MBD, one can clearly see that there are significant differences compared to PBE+MBD. It can be seen that the energy barrier is with 4 kJ/mol much smaller and the energy difference between the two polymorphs is now larger than 3 kJ/mol. When we utilize the DFTB3+MBD structures and perform on top of them PBE+MBD/light single point calculations, we get similar results to the PBE+MBD NEB results, but the maximum is slightly shifted. This suggests that DFTB+MBD is not reliable in this context when it comes to energies, but the structures along the minimum energy path seem to have a better quality.
Our SO3krates MLFF performed very well, the differences to the PBE+MBD NEB results and subsequent PBE+MBD single point calculations are less than 0.5 kJ/mol. When we analyze the differences between the SO3krates results and the PBE+MBD single points on the same structures along the minimum energy path, we obtain a mean absolute error (MAE) and mean error (ME) for the relative energies of only 0.2 and 0.1 kJ/mol, respectively. Performing the NEB calculations with 15 instead of the otherwise used 10 images, yields virtually identical results (within 0.1 kJ/mol). Hence, our MLFF yields very accurate results and by far outperformed DFTB3+MBD. Therefore, we have performed the full NEB calculations for the remaining two systems only with the generated MLFF.
Since it is possible for this system to utilize a simple linear interpolation as starting point of NEB, we have compared the efficiency with our new interpolation method. Both methods yield identical results, but our new methodology converged slightly faster in 182 steps compared to 206 steps of the simple linear interpolation.
After having discussed the energetics, we analyze the obtained transition-state geometries. The unit cell of the PBE+MBD transition state is visualized in Fig. 1. It can be seen that the orientation of the left two molecules is exactly in between both polymorphs. To assess the quality of our PBE+MBD transition-state geometry, we have calculated the phonon frequencies and find two imaginary frequencies at the -point: -73.6 cm-1, which corresponds to the movement along the NEB pathway, and another mode at -14.5 cm-1. At a true transition state, we should only observe the first imaginary mode. Therefore, we have slightly displaced the structure along the second imaginary mode, followed by a subsequent phonon calculation. In this way we were able to fully remove this second imaginary mode and have obtained a true transition state. The differences between the two structures are essentially negligible, with a lattice energy difference of 0.014 kJ/mol and a RMSD20 value of 0.038 Å, which is the RMSD value of a cluster consisting of 20 molecules. When we compare the obtained transition states of DFTB3+MBD and SO3krates with the PBE+MBD results, we obtain RMSD20 values of 0.308 Å and 0.193 Å, respectively. Also in this case, the MLFF outperforms DFTB3+MBD and yields a structure very close to the PBE+MBD result.
III.2 Theophylline Polymorph Transition
Next, we discuss the polymorphic transition from theophylline polymorph I Khamar et al. (2011), which is stable at higher temperatures, to the room-temperature stable polymorph II Fucke et al. (2012b). This transition (I to II) does not occur upon cooling due to significant structural differences.Fucke et al. (2012a) However, the reverse transition of polymorph II into I is known to occur upon heating. Since we can expect here a lot of atomic motion, this system is ideal for testing our new interpolation method. When dividing every molecule into three fragments (see Fig. 2), our new interpolation method automatically creates initial reaction path structures without any atomic collisions.
The results of the minimum-energy path obtained with the SO3krates MLFF together with single point calculations on top of the SO3krates structures are shown in Fig. 4. In both cases, polymorph II is, as expected, thermodynamically more stable by about 5 kJ/mol. In both cases we obtain a substantial energy barrier, amounting to 46.7 kJ/mol and 49.7 kJ/mol for SO3krates and the PBE+MBD single points, respectively. SO3krates provides again very accurate results, underestimating the PBE+MBD energy barrier by only 6 % and with a MAE and ME across the minimum energy path of 0.9 and 0.8 kJ/mol, respectively. The obtained transition state is shown in Fig. 1 and one can see that this process involves substantial motion of all molecules, given the large changes in unit-cell parameters.
III.3 Solid-State Diels-Alder Reaction
Finally, we investigate if our approach can also be applied to solid-state reactions. As example we utilize a single-crystal to single-crystal Diels-Alder reaction of a co-crystal composed of bis(N-allylimino)-1,4-dithiin and 9-bromoanthracene.Khorasani et al. (2015) Experimentally, it took 84 days to achieve a 97 % conversion.
When dividing both molecules building up the co-crystal into several fragments (see Fig. 2), our interpolation yielded also for this reaction automatically collision-free initial path structures. Fig. 5 shows again the results of the minimum-energy path obtained with the SO3krates MLFF together with single point calculations on top of the SO3krates structures. In both cases, the product of the reaction is about 5.5 kJ/mol more stable than the co-crystal formed by the reactants. The energy barrier is for both methods very similar amounting for SO3krates and PBE+MBD to 27.7 and 28.6 kJ/mol, respectively. The deviation of the MLFF energy barrier compared to the DFT reference is less than 1 kJ/mol and statistically over the whole minimum energy path the MAE and ME amount to only 0.5 and -0.1 kJ/mol, respectively. The corresponding transition state is visualized in Fig. 1. Before the actual ring closure can happen, several propenyl groups need to re-arrange themselves, which apparently is the main source of the energy barrier in this case.
IV Conclusion
We have introduced a novel interpolation method for generating an initial pathway for subsequent nudged-elastic band (NEB) calculations suitable for periodic structures like molecular crystals. This methodology aims to automatically avoid any atomic collisions, which otherwise regularly happen in periodic systems when using simple linear interpolations. We demonstrated the versatility of this approach by successfully applying it to a range of divers systems, including two polymorph transitions — one with a small and one with a rather large energy barrier — as well as a single-crystal Diels-Alder reaction. Our approach consistently provided collision-free initial pathways, avoiding the need for cumbersome manual corrections. This automated pathway generation greatly simplifies the initial setup for NEB calculations, making it a highly practical tool for tackling even complex periodic systems.
Furthermore, we have utilized the SO3krates architecture to train an individual machine-learned force field (MLFF) for every studied system. By training the MLFFs on a diverse set of reference data obtained from PBE+MBD calculations, we were able to reproduce the energy barriers and structural characteristics of the minimum energy pathways with high accuracy. The MAEs between the MLFF energies along the reaction paths compared to subsequent PBE+MBD single-point energies remained consistently below 1 kJ/mol for all studied systems. However, we note that the reference data focused heavily on the PES region required for the studied transition and the resulting MLFFs will likely be much less accurate when simulating for instance other polymorphs.
Finally, many polymorph transition and reaction paths can probably not be described by utilizing unit cells but would require larger supercells to allow for the needed flexibility. In such cases, computationally efficient methods like MLFFs are key for being able to afford the necessary calculations. Our results highlight that with properly trained MLFFs, it is possible to achieve both computational efficiency and accuracy, making them a very promising tool for exploring complex solid-state phenomena.
Acknowledgements.
The computational results presented have been achieved in part using the Vienna Scientific Cluster (VSC).References
- Kaupp (2002) G. Kaupp, Curr. Opin. Solid State Mater. Sci. 6, 131–138 (2002).
- Datta and Grant (2004) S. Datta and D. J. W. Grant, Nat. Rev. Drug Discovery 3, 42 (2004).
- Anthony (2006) J. E. Anthony, Chem. Rev. 106, 5028–5048 (2006).
- Saha, Nath, and Thakuria (2022) B. K. Saha, N. K. Nath, and R. Thakuria, Chem. Rec. 23, e202200173 (2022).
- Zhou et al. (2018) Y. Zhou, J. Wang, Y. Xiao, T. Wang, and X. Huang, Curr. Pharm. Des. 24, 2375–2382 (2018).
- Galwey (2014) A. K. Galwey, React. Kinet. Mech. Catal. 114, 1–29 (2014).
- Jónsson, Mills, and Jacobsen (1998) H. Jónsson, G. Mills, and K. W. Jacobsen, in Classical and Quantum Dynamics in Condensed Phase Simulations (World Scientific 1998, 1998) p. 385–404.
- Henkelman, Uberuaga, and Jónsson (2000) G. Henkelman, B. P. Uberuaga, and H. Jónsson, J. Chem. Phys. 113, 9901 (2000).
- Sheppard et al. (2012) D. Sheppard, P. Xiao, W. Chemelewski, D. D. Johnson, and G. Henkelman, J. Chem. Phys. 136, 074103 (2012).
- Mader et al. (2018) S. L. Mader, A. Bräuer, M. Groll, and V. R. I. Kaila, Nat. Commun. 9, 1168 (2018).
- Shin et al. (2019) J. Shin, D. S. Choi, H. J. Lee, Y. Jung, and J. W. Choi, Adv. Energy Mater. 9, 1900083 (2019).
- Han and Luber (2020) R. Han and S. Luber, J. Comput. Chem. 41, 1586–1597 (2020).
- Hossain et al. (2020) M. D. Hossain, Y. Huang, T. H. Yu, W. A. Goddard III, and Z. Luo, Nat. Commun. 11, 2256 (2020).
- Podewitz, Sen, and Buchmeiser (2021) M. Podewitz, S. Sen, and M. R. Buchmeiser, Organometallics 40, 2478–2488 (2021).
- Salvalaglio et al. (2014) M. Salvalaglio, C. Perego, F. Giberti, M. Mazzotti, and M. Parrinello, Proc. Natl. Acad. Sci. U.S.A 112, E6 (2014).
- Gimondi and Salvalaglio (2017) I. Gimondi and M. Salvalaglio, J. Chem. Phys. 147, 114502 (2017).
- Francia, Price, and Salvalaglio (2021) N. F. Francia, L. S. Price, and M. Salvalaglio, CrystEngComm 23, 5575–5584 (2021).
- Otero-de-la-Roza and Johnson (2012) A. Otero-de-la-Roza and E. R. Johnson, J. Chem. Phys. 137, 054103 (2012).
- Reilly and Tkatchenko (2013) A. M. Reilly and A. Tkatchenko, J. Chem. Phys. 139, 024705 (2013).
- Marom et al. (2013) N. Marom, R. A. DiStasio, V. Atalla, S. Levchenko, A. M. Reilly, J. R. Chelikowsky, L. Leiserowitz, and A. Tkatchenko, Angew. Chem. Int. Ed. 52, 6629–6632 (2013).
- Reilly and Tkatchenko (2015) A. M. Reilly and A. Tkatchenko, Chem. Sci. 6, 3289 (2015).
- Brandenburg and Grimme (2016) J. G. Brandenburg and S. Grimme, Acta Crystallogr., Sect. B: Struct. Sci. 72, 502–513 (2016).
- Shtukenberg et al. (2017) A. G. Shtukenberg, Q. Zhu, D. J. Carter, L. Vogt, J. Hoja, E. Schneider, H. Song, B. Pokroy, I. Polishchuk, A. Tkatchenko, A. R. Oganov, A. L. Rohl, M. E. Tuckerman, and B. Kahr, Chem. Sci. 8, 4926 (2017).
- Hoja et al. (2019) J. Hoja, H.-Y. Ko, M. A. Neumann, R. Car, R. A. DiStasio, and A. Tkatchenko, Sci. Adv. 5, eaau3338 (2019).
- Dolgonos, Hoja, and Boese (2019) G. A. Dolgonos, J. Hoja, and A. D. Boese, Phys. Chem. Chem. Phys. 21, 24333 (2019).
- Jana et al. (2021) S. Jana, H. Myneni, S. Śmiga, L. A. Constantin, and P. Samal, J. Chem. Phys. 155, 114102 (2021).
- Price, Otero-de-la-Roza, and Johnson (2023) A. J. A. Price, A. Otero-de-la-Roza, and E. R. Johnson, Chem. Sci. 14, 1252 (2023).
- Hoja, List, and Boese (2023) J. Hoja, A. List, and A. D. Boese, J. Chem. Theory Comput. 20, 357–367 (2023).
- Schütt et al. (2017) K. T. Schütt, F. Arbabzadah, S. Chmiela, K. R. Müller, and A. Tkatchenko, Nat. Commun. 8, 13890 (2017).
- Schütt, Unke, and Gastegger (2021) K. Schütt, O. Unke, and M. Gastegger, in Proceedings of the 38th International Conference on Machine Learning, Proceedings of Machine Learning Research, Vol. 139, edited by M. Meila and T. Zhang (PMLR, 2021) pp. 9377–9388.
- Unke et al. (2021) O. T. Unke, S. Chmiela, M. Gastegger, K. T. Schütt, H. E. Sauceda, and K.-R. Müller, Nat. Commun. 12, 7273 (2021).
- Batzner et al. (2022) S. Batzner, A. Musaelian, L. Sun, M. Geiger, J. P. Mailoa, M. Kornbluth, N. Molinari, T. E. Smidt, and B. Kozinsky, Nat. Commun. 13, 2453 (2022).
- Batatia et al. (2022) I. Batatia, D. P. Kovacs, G. Simm, C. Ortner, and G. Csanyi, in Advances in Neural Information Processing Systems, Vol. 35, edited by S. Koyejo, S. Mohamed, A. Agarwal, D. Belgrave, K. Cho, and A. Oh (Curran Associates, Inc., 2022) pp. 11423–11436.
- Frank, Unke, and Müller (2022) T. Frank, O. Unke, and K.-R. Müller, Advances in Neural Information Processing Systems 35, 29400 (2022).
- Frank et al. (2024) T. Frank, O. Unke, K.-R. Müller, and S. Chmiela, Nat. Commun. 15, 6539 (2024).
- Alcobé et al. (1994) X. Alcobé, E. Estop, A. E. Aliev, K. D. Harris, J. Rodriguez-Carvajal, and J. Rius, J. Solid State Chem. 110, 20 (1994).
- Khamar et al. (2011) D. Khamar, R. G. Pritchard, I. J. Bradshaw, G. A. Hutcheon, and L. Seton, Acta Crystallogr., Sect. C: Cryst. Struct. Commun. 67, 496–499 (2011).
- Fucke et al. (2012a) K. Fucke, G. J. McIntyre, C. Wilkinson, M. Henry, J. A. K. Howard, and J. W. Steed, Cryst. Growth Des. 12, 1395–1401 (2012a).
- Khorasani et al. (2015) S. Khorasani, D. S. Botes, M. A. Fernandes, and D. C. Levendis, CrystEngComm 17, 8933–8945 (2015).
- Perdew, Burke, and Ernzerhof (1996) J. P. Perdew, K. Burke, and M. Ernzerhof, Phys. Rev. Lett. 77, 3865 (1996).
- Tkatchenko et al. (2012) A. Tkatchenko, R. A. DiStasio, R. Car, and M. Scheffler, Phys. Rev. Lett. 108, 236402 (2012).
- Ambrosetti et al. (2014) A. Ambrosetti, A. M. Reilly, R. A. DiStasio, and A. Tkatchenko, J. Chem. Phys. 140, 18A508 (2014).
- Hermann and Tkatchenko (2020) J. Hermann and A. Tkatchenko, Phys. Rev. Lett. 124, 146401 (2020).
- Blum et al. (2009) V. Blum, R. Gehrke, F. Hanke, P. Havu, V. Havu, X. Ren, K. Reuter, and M. Scheffler, Comput. Phys. Commun. 180, 2175 (2009).
- Knuth et al. (2015) F. Knuth, C. Carbogno, V. Atalla, V. Blum, and M. Scheffler, Comput. Phys. Commun. 190, 33 (2015).
- Ren et al. (2012) X. Ren, P. Rinke, V. Blum, J. Wieferink, A. Tkatchenko, A. Sanfilippo, K. Reuter, and M. Scheffler, New J. Phys. 14, 053020 (2012).
- Yu et al. (2018) V. W. Yu, F. Corsetti, A. García, W. P. Huhn, M. Jacquelin, W. Jia, B. Lange, L. Lin, J. Lu, W. Mi, A. Seifitokaldani, Á. Vázquez-Mayagoitia, C. Yang, H. Yang, and V. Blum, Comput. Phys. Commun. 222, 267 (2018).
- Havu et al. (2009) V. Havu, V. Blum, P. Havu, and M. Scheffler, J. Comput. Phys. 228, 8367 (2009).
- Ihrig et al. (2015) A. C. Ihrig, J. Wieferink, I. Y. Zhang, M. Ropo, X. Ren, P. Rinke, M. Scheffler, and V. Blum, New J. Phys. 17, 093020 (2015).
- Larsen et al. (2017) A. H. Larsen, J. J. Mortensen, J. Blomqvist, I. E. Castelli, R. Christensen, M. Dułak, J. Friis, M. N. Groves, B. Hammer, C. Hargus, E. D. Hermes, P. C. Jennings, P. B. Jensen, J. Kermode, J. R. Kitchin, E. L. Kolsbjerg, J. Kubal, K. Kaasbjerg, S. Lysgaard, J. B. Maronsson, T. Maxson, T. Olsen, L. Pastewka, A. Peterson, C. Rostgaard, J. Schiøtz, O. Schütt, M. Strange, K. S. Thygesen, T. Vegge, L. Vilhelmsen, M. Walter, Z. Zeng, and K. W. Jacobsen, J. Phys.: Condens.Matter 29, 273002 (2017).
- Shoemake (1985) K. Shoemake, in Proceedings of the 12th annual conference on Computer graphics and interactive techniques, SIGGRAPH 85 (ACM Press, 1985) p. 245–254.
- Nguyen, Jaillet, and Redon (2017) M. K. Nguyen, L. Jaillet, and S. Redon, J. Comput.-Aided Mol. Des. 31, 403–417 (2017).
- Nguyen, Jaillet, and Redon (2018) M. K. Nguyen, L. Jaillet, and S. Redon, J. Comput.-Aided Mol. Des. 32, 853–867 (2018).
- Togo and Tanaka (2015) A. Togo and I. Tanaka, Scr. Mater. 108, 1 (2015).
- Hourahine et al. (2020) B. Hourahine, B. Aradi, V. Blum, F. Bonafé, A. Buccheri, C. Camacho, C. Cevallos, M. Y. Deshaye, T. Dumitrică, A. Dominguez, S. Ehlert, M. Elstner, T. van der Heide, J. Hermann, S. Irle, J. J. Kranz, C. Köhler, T. Kowalczyk, T. Kubař, I. S. Lee, V. Lutsker, R. J. Maurer, S. K. Min, I. Mitchell, C. Negre, T. A. Niehaus, A. M. N. Niklasson, A. J. Page, A. Pecchia, G. Penazzi, M. P. Persson, J. Řezáč, C. G. Sánchez, M. Sternberg, M. Stöhr, F. Stuckenberg, A. Tkatchenko, V. W.-z. Yu, and T. Frauenheim, J. Chem. Phys. 152, 124101 (2020).
- Langer, Frank, and Knoop (2023) M. F. Langer, J. T. Frank, and F. Knoop, J. Chem. Phys. 159, 174105 (2023).
- Fucke et al. (2012b) K. Fucke, G. McIntyre, C. Wilkinson, M. Henry, J. Howard, and J. Steed, “Ccdc 878133: Experimental crystal structure determination,” (2012b).